OpenStudy (anonymous):

integrate 2 / x(x^2+1)^2

OpenStudy (anonymous):

try substitution t=x^2 dt=2xdx 1/2dt=xdx

OpenStudy (anonymous):

can i do partial fraction method?

OpenStudy (agent0smith):

\[\int\limits \frac{ 2 }{ x (x^2+1)^2} dx\] Hmm. Use partial fractions?

OpenStudy (anonymous):

yes you can use partial fraction . i guess fractions will be easy with linear terms so use t=x^2 dt=2xdx \[\Large \int\limits \frac{dt}{t(t+1)^2}\] now use partial fractions .

OpenStudy (anonymous):

no we can write the original integral as \[\Large \int\limits \frac{2xdx}{x^2(x^2+1)}\] so let t=x^2 dt=2xdx

OpenStudy (anonymous):

\[2\left( \frac{ 1 }{ 2(x^2+1) }-\frac{ 1 }{ 2 }\log(x^2+1)+\log x \right)\]

OpenStudy (anonymous):

@sami-21 i dont understand

OpenStudy (anonymous):

you can use partial fraction, this is because if the power of the denominator is greater than the power of the numerator that would be an alternative method...

OpenStudy (anonymous):

but for me personal i prefer the first method present by sami....

OpenStudy (anonymous):

okay thank you!

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

did u solve the problem?